OpenStudy (anonymous):

determine the following identity: 1 = cos2A on cos2A equals : tan2A on tanA

OpenStudy (anonymous):

Substitute \[\cos2A = {2\tan A \over 1+\tan^2 A}\] and you will get your answer

OpenStudy (anonymous):

Is this your question \[{1- \cos 2A \over \cos 2A}={\tan 2A \over \tan A}\]

OpenStudy (anonymous):

@Luahn ??

OpenStudy (anonymous):

yes thats the question please

OpenStudy (anonymous):

but its + cos2A

OpenStudy (anonymous):

1+ co2A

OpenStudy (anonymous):

Ok. just write it as \[{1+ \cos 2A \over \cos 2A}\] Now substitute instead of cos 2A, \[\cos2A = {2\tan A \over 1+\tan^2 A}\] and simplify

OpenStudy (anonymous):

i need to prove that 1+cos2A cos 2A equals: tan2A tanA

OpenStudy (anonymous):

can i please have the method of getting to the answer